# An Unsupervised Approach to Achieve Supervised-Level Explainability in Healthcare Records

### v1

@@ 1 Introduction
Feature attributions, a common form of explainability, can help healthcare professionals quickly find the evidence necessary to review a medical code suggestion (see fig. 1). Feature attribution methods score each input feature based on its influence on the model’s output. These explanations are often evaluated through plausibility and faithfulness Jacovi and Goldberg (2020). Plausibility measures how convincing an explanation is to human users, while faithfulness measures the explanation’s ability to reflect the model’s logic.
@@ 2.2 Adversarial robustness and explainability
Adversarial robustness refers to the ability of a machine learning model to maintain performance under adversarial attacks, which involve making small changes to input data that do not significantly affect human perception or judgment (e.g., a small amount of image noise). Tsipras et al. (2019) and Ilyas et al. (2019) demonstrate that adversarial examples exploit the models’ dependence on fragile, non-robust features. Adversarial robustness training embeds invariances to prevent models relying on such non-robust features, with regularization and data augmentation as main strategies Tsipras et al. (2019); Ross and Doshi-Velez (2017).
Previous work in image classification shows that adversarially robust models generate more plausible explanations Ross and Doshi-Velez (2017); Chen et al. (2019); Etmann et al. (2019). These studies demonstrate this phenomenon for three adversarial training strategies: 1) input gradient regularization, which improves the Lipschitzness of neural networks Drucker and Le Cun (1992); Ross and Doshi-Velez (2017); Chen et al. (2019); Etmann et al. (2019); Rosca et al. (2020); Fel et al. (2022); Khan et al. (2023), 2) adversarial training, which trains models on adversarial examples, thereby embeds invariance to adversarial noise Tsipras et al. (2019), and 3) feature masking, which masks unimportant features during training to embed invariance to such features Bhalla et al. (2023).
The relationship between robustness and explanation plausibility in NLP is unclear, as tokens differ from pixels. To our knowledge, Yoo and Qi (2021) and Li et al. (2023) are the only studies investigating this. However, these studies evaluate the faithfulness of explanations using only the simple task of sentiment classification.
@@ 3.2 Feature attribution methods
We evaluated several feature attribution methods for automated medical coding, categorizing them into three types: attention-based, gradient-based, and perturbation-based (more details in appendix B). Attention-based methods like Attention Mullenbach et al. (2018), Attention Rollout Abnar and Zuidema (2020), and AttGrad Serrano and Smith (2019) rely on the model’s attention weights. Gradient-based methods such as InputXGrad Sundararajan et al. (2017), Integrated Gradients Sundararajan et al. (2017), and Deeplift Shrikumar et al. (2017) use backpropagation to quantify the influence of input features on outputs. Perturbation-based methods, including LIME Ribeiro et al. (2016), KernelSHAP Lundberg and Lee (2017), and Occlusion@1 Ribeiro et al. (2016), measure the impact on output confidence by occluding input features.
Our preliminary analysis showed that Attention and InputXGrad often attributed high importance to non-informative tokens. However, the two methods rarely attributed high importance to the same irrelevant tokens. Therefore, we propose a new feature attribution method, AttInGrad, combining Attention and InputXGrad by multiplying their attribution scores, thereby down-prioritizing irrelevant tokens. We calculated the AttInGrad attribution scores for class using the following equation:
@@ 4.2 Models
We used PLM-ICD, a state-of-the-art automated medical coding model architecture, for our experiments because its architecture is simple while outperforming other models according to Huang et al. (2022); Edin et al. (2023). To address stability issues caused by numerical overflow in the decoder of the original model, we replaced the label-wise attention mechanism with standard cross-attention Vaswani et al. . This adjustment not only stabilized training but also slightly improved performance. We provide further details on the architecture modifications in appendix A.
@@ 5 Results
The objective of this paper was to produce high-quality explanations without relying on evidence span annotations. In fig. 2, we compare the plausibility of our approach (Token masking and AttnInGrad) with the unsupervised approach (B and Attention) and supervised state-of-the-art approach (B and Attention). Our approach was substantially more plausible than the unsupervised on all metrics. Compared with the supervised, our approach achieved similar F1 and Recall@5 and substantially better Empty scores. The supervised approach achieved similar plausibility to ours on most metrics (see table 2). Our approach also achieved the highest comprehensiveness and lowest sufficiency scores (see fig. 3). The difference was larger in the sufficiency scores, where the supervised score was twice as high as ours.
@@
We evaluated the explanation plausibility of every model and explanation method combination in table 2. IGR and TM outperformed the baseline model on most metrics and explanation methods. In section E.5, we compare the unsupervised models on a bigger test set and see similar results. The supervised model yielded better results for attention-based explanations but was weaker than the robust models when using the gradient-based explanation methods: , IG, and Deeplift.
@@
AttInGrad was also more faithful than Attention (see fig. 3). However, while AttInGrad surpassed the gradient-based methods in comprehensiveness, its sufficiency scores were slightly worse.
@@
In fig. 4, we analyze the relationship between explanation quality (y-axis) and the proportion of the top five most important tokens that are special tokens (x-axis). Each data point represents the average statistics across the MDACE test set for one seed/run of the B model. Figures 4(a) and 4(b) show F1 (plausibility) and comprehensiveness (faithfulness) respectively. For Attention and AttInGrad, we see strong negative correlations for both metrics with a large inter-seed variance. The regression lines fitted on Attention and AttInGrad overlap, with the data points from AttInGrad shifted slightly towards the upper left, indicating attribution of less importance to special tokens.
@@ 6 Discussion
Intuitively, training a model on evidence spans should encourage it to use features relevant to humans, thereby making its explanations more plausible. However, we hypothesize that the training strategy used by Cheng et al. (2023) primarily addresses the shortcomings of attention-based explanation methods rather than enhancing the model’s underlying logic. The model B only produced more plausible explanations with attention-based feature attribution methods (see fig. 2). If the model truly leveraged more informative features, we would expect to see improvements across various feature attribution methods. Additionally, the differences between Attention and AttInGrad were negligible for B compared to the other models. This may suggest that the supervised training might have corrected some of the inherent issues in the Attention method, similar to what AttInGrad achieves.
@@ 7 Conclusion
Our study did not conclusively demonstrate why adversarial robustness training strategies improved the explanation plausibility. We hypothesized that these strategies force the model to rely on fewer features that weakly correlate with the labels, and such features are less plausible. However, validating this hypothesis proved challenging. Our analysis of feature attributions’ entropy was inconclusive, as detailed in section E.4. Moreover, we did not know which features the model relied on because this would require a perfect feature attribution method, which is what we aimed to develop. Despite these challenges, we demonstrated that the adversarial robust models produced more plausible explanations. We believe that our work has laid a solid foundation for future research into how model training strategies can impact explanation plausibility.
@@ Acknowledgements
This research was partially funded by the Innovation Fund Denmark via the Industrial Ph.D. Program (grant no. 2050-00040B) and Academy of Finland (grant no. 322653). We thank Jonas Lyngsø for insightful discussions and code for dataloading. Furthermore, we thank Simon Flachs, Lana Krumm, and Andreas Geert Motzfeldt for revisions.
@@ Appendix A Model architecture details
PLM-ICD is a state-of-the-art automated medical coding model Huang et al. (2022); Edin et al. (2023). We experienced that PLM-ICD occasionally crashed during training. Therefore, we modified the architecture and called it pre-trained language model with class-wise cross attention (PLM-CA) Huang et al. (2022). Our architecture comprises an encoder and a decoder (see fig. 5). The encoder transforms a sequence of tokens indices into a sequence of contextualized token representations . Both PLM-ICD and PLM-CA use RoBERTa-PM, a transformer pre-trained on PubMed articles and clinical notes, as the encoder Lewis et al. (2020).
@@ Appendix B Feature attribution methods
Attention ()
We use the raw attention weights (see eq. 2) in the cross-attention layer to explain class . As mentioned in section 2.1, this explanation method was used by most previous studies in automated medical coding Mullenbach et al. (2018); Kim et al. (2022); Dong et al. (2021); Teng et al. (2020); Cheng et al. (2023).
@@
The attention matrix in the cross-attention layer extracts information from the contextualized token representations encoded by RoBERTa (see fig. 5). The token representations are not guaranteed to be aligned with the input tokens. A token representation at position could represent any and multiple tokens in the document. Attention rollout considers all the model’s attention layers to calculate the feature attributions Abnar and Zuidema (2020). First, the attention matrices in each layer are averages across the heads. Then, the identity matrix is added to each layer’s attention matrix to represent the skip connections. Finally, the attention rollout is calculated recursively using eq. 4.
@@
Integrated Gradients (IG)
Integrated Gradients (IG) assigns an attribution score to each input feature by computing the integral of the gradients along the straight line path from the baseline to the input Sundararajan et al. (2017). Similar to InputXGradient, we used the L2-norm of the output to get the final attribution scores.
@@
AttnGrad ()
AttnGrad multiplies the attention with the gradient of the model’s output with respect to the attention weights Serrano and Smith (2019):
@@ E.2 Results from all feature attribution methods
In the main paper, we only presented the results of selected feature attribution methods because of space constraints. Here, we present the results for all the feature attribution methods: Attention (), AttGrad (), Attention Rollout (Rollout), InputXGrad (), Integrated gradients (IG), Deeplift, and AttInGrad (). We compare these methods with a random baseline (Rand), which randomly generates attribution scores. We present the plausibility results in table 7, and the faithfulness results in table 8.
@@ E.5 Unsupervised comparison on bigger test set
We present the plausibility results in table 12. We observe that the results are similar to those of the main paper. However, the IGR produced substantially better attention-based explanations than in the main paper. In fig. 6, we inspect the inter-seed variance. We observe that IGR has no outliers. We, therefore, attribute the differences between this comparison and that in the main paper to none of the ten IGR runs happening to produce an outlier model. These results highlight the fragility of evaluating the attention-based feature attribution methods.

### v2

@@ 1 Introduction
Feature attributions, a common form of explainability, can help healthcare professionals quickly find the evidence necessary to review a medical code suggestion (see Figure 1). Feature attribution methods score each input feature based on its influence on the model’s output. These explanations are often evaluated through plausibility and faithfulness Jacovi and Goldberg (2020). Plausibility measures how convincing an explanation is to human users, while faithfulness measures the explanation’s ability to reflect the model’s logic.
@@ 2.2 Adversarial robustness and explainability
Adversarial robustness refers to the ability of a machine learning model to maintain performance under adversarial attacks, which involve making small changes to input data that do not significantly affect human perception or judgment (e.g., a small amount of image noise). Tsipras et al. (2018) and Ilyas et al. (2019) demonstrate that adversarial examples exploit the models’ dependence on fragile, non-robust features. Adversarial robustness training embeds invariances to prevent models relying on such non-robust features, with regularization and data augmentation as main strategies Tsipras et al. (2018); Ros and Doshi-Velez (2018).
Previous work in image classification shows that adversarially robust models generate more plausible explanations Ros and Doshi-Velez (2018); Chen et al. (2019); Etmann et al. (2019). These studies demonstrate this phenomenon for three adversarial training strategies: 1) input gradient regularization, which improves the Lipschitzness of neural networks Drucker and Le Cun (1992); Ros and Doshi-Velez (2018); Chen et al. (2019); Etmann et al. (2019); Rosca et al. (2020); Fel et al. (2022); Khan et al. (2023), 2) adversarial training, which trains models on adversarial examples, thereby embeds invariance to adversarial noise Tsipras et al. (2018), and 3) feature masking, which masks unimportant features during training to embed invariance to such features Bhalla et al. (2023).
The relationship between model robustness and explanation plausibility in Natural Language Processing (NLP) remains unclear, primarily due to the fundamental differences between textual tokens and image pixels. To date, only two studies, Yoo and Qi (2021) and Li et al. (2023), have explored this relationship in depth. These studies suggest that robust models produce more faithful explanations compared to their non-robust counterparts. However, their conclusions may be questionable due to the use of the Area Over the Perturbation Curve (AOPC) metric, which is unsuitable for cross-model comparisons Edin et al. (2024). This inappropriate application of AOPC may have led to potentially misleading conclusions.
@@ 3.2 Feature attribution methods
We evaluated several feature attribution methods for automated medical coding, categorizing them into three types: attention-based, gradient-based, and perturbation-based (more details in Appendix B). Attention-based methods like Attention Mullenbach et al. (2018), Attention Rollout Abnar and Zuidema (2020), and AttGrad Serrano and Smith (2019) rely on the model’s attention weights. Gradient-based methods such as InputXGrad Sundararajan et al. (2017), Integrated Gradients (IntGrad) Sundararajan et al. (2017), and Deeplift Shrikumar et al. (2017) use backpropagation to quantify the influence of input features on outputs. Perturbation-based methods, including LIME Ribeiro et al. (2016), KernelSHAP Lundberg and Lee (2017), and Occlusion@1 Ribeiro et al. (2016), measure the impact on output confidence by occluding input features.
Our investigation into feature attribution methods revealed an intriguing pattern: while individual methods often produced unreliable explanations, their shortcomings rarely overlapped. Attention-based methods and gradient-based approaches like InputXGrad frequently disagreed on which tokens were most important, yet both contributed valuable insights in different scenarios. This observation sparked a key question: could we leverage the complementary strengths of these methods to create a more robust attribution technique?
To address this question, we propose AttInGrad, a novel feature attribution method that combines Attention and InputXGrad. AttInGrad multiplies their respective attribution scores, aiming to amplify the importance of tokens deemed relevant by both methods while down-weighting those highlighted by only one or neither method.
We formalize the AttInGrad attribution scores for class using the following equation:
@@
In Section 5, we will provide an in-depth analysis of the mechanisms underlying AttInGrad’s effectiveness, shedding light on why this combination of methods yields improved feature attributions.
@@ 4.2 Models
We used PLM-ICD, a state-of-the-art automated medical coding model architecture, for our experiments because its architecture is simple while outperforming other models according to Huang et al. (2022); Edin et al. (2023). To address stability issues caused by numerical overflow in the decoder of the original model, we replaced the label-wise attention mechanism with standard cross-attention Vaswani et al. (2017). This adjustment not only stabilized training but also slightly improved performance. We provide further details on the architecture modifications in Appendix A.
@@ 5 Results
The objective of this paper was to produce high-quality explanations without relying on evidence span annotations. In Figure 2, we compare the plausibility of our approach (Token masking and AttnInGrad) with the unsupervised approach (B and Attention) and supervised state-of-the-art approach (B and Attention). Our approach was substantially more plausible than the unsupervised on all metrics. Compared with the supervised, our approach achieved similar F1 and Recall@5 and substantially better Empty scores. The supervised approach achieved similar plausibility to ours on most metrics (see Table 2). Our approach also achieved the highest comprehensiveness and lowest sufficiency scores (see Figure 3). The difference was larger in the sufficiency scores, where the supervised score was twice as high as ours.
@@
We evaluated the explanation plausibility of every model and explanation method combination in Table 2. IGR and TM outperformed the baseline model on most metrics and explanation methods. In Section E.5, we compare the unsupervised models on a bigger test set and see similar results. The supervised model yielded better results for attention-based explanations but was weaker than the robust models when using the gradient-based explanation methods: InputXGrad, IG, and Deeplift.
@@
AttInGrad was also more faithful than Attention (see Figure 3). However, while AttInGrad surpassed the gradient-based methods in comprehensiveness, its sufficiency scores were slightly worse.
@@
In Figure 4, we analyze the relationship between explanation quality (y-axis) and the proportion of the top five most important tokens that are special tokens (x-axis). Each data point represents the average statistics across the MDACE test set for one seed/run of the B model. Figures 4(a) and 4(b) show F1 (plausibility) and comprehensiveness (faithfulness) respectively. For Attention and AttInGrad, we see strong negative correlations for both metrics with a large inter-seed variance. The regression lines fitted on Attention and AttInGrad overlap, with the data points from AttInGrad shifted slightly towards the upper left, indicating attribution of less importance to special tokens.
@@ 6 Discussion
Intuitively, training a model on evidence spans should encourage it to use features relevant to humans, thereby making its explanations more plausible. However, we hypothesize that the training strategy used by Cheng et al. (2023) primarily addresses the shortcomings of attention-based explanation methods rather than enhancing the model’s underlying logic. The model B only produced more plausible explanations with attention-based feature attribution methods (see Figure 2). If the model truly leveraged more informative features, we would expect to see improvements across various feature attribution methods. Additionally, the differences between Attention and AttInGrad were negligible for B compared to the other models. This may suggest that the supervised training might have corrected some of the inherent issues in the Attention method, similar to what AttInGrad achieves.
@@ 7 Conclusion
Our study did not conclusively show why adversarial robustness training strategies improved the explanation plausibility. We hypothesized that these strategies force the model to rely on fewer features that weakly correlate with the labels, and such features are less plausible. However, validating this hypothesis proved challenging. Our analysis of feature attributions’ entropy was inconclusive, as detailed in Section E.4. Moreover, we did not know which features the model relied on because this would require a perfect feature attribution method, which is what we aimed to develop. Despite these challenges, we demonstrated that the adversarial robust models produced more plausible explanations. We believe that our work has laid a solid foundation for future research into how model training strategies can impact explanation plausibility.
Our study’s scope was constrained to a single data source (Beth Israel Deaconess Medical Center’s MIMIC-III and MDACE) and one model architecture (PLM-ICD). The effectiveness of Attention and AttInGrad may vary with different architectures, particularly those employing multi-head attention and skip connections in the final layer. Further research is needed to investigate the generalizability of our findings across diverse model architectures, medical coding systems, languages, and healthcare institutions.
@@ Acknowledgements
This research was partially funded by the Innovation Fund Denmark via the Industrial Ph.D. Program (grant no. 2050-00040B) and Research Council of Finland (grant no. 322653). We thank Jonas Lyngsø for insightful discussions and code for dataloading. Furthermore, we thank Simon Flachs, Lana Krumm, and Andreas Geert Motzfeldt for revisions.
@@ Appendix A Model architecture details
PLM-ICD is a state-of-the-art automated medical coding model Huang et al. (2022); Edin et al. (2023). It comprises 131 million parameters. We experienced that PLM-ICD occasionally crashed during training. Therefore, we modified the architecture and called it pre-trained language model with class-wise cross attention (PLM-CA) Huang et al. (2022). Our architecture comprises an encoder and a decoder (see Figure 5). The encoder transforms a sequence of tokens indices into a sequence of contextualized token representations . Both PLM-ICD and PLM-CA use RoBERTa-PM, a transformer pre-trained on PubMed articles and clinical notes, as the encoder Lewis et al. (2020).
@@ Appendix B Feature attribution methods
Attention
We use the raw attention weights (see Equation 2) in the cross-attention layer to explain class . As mentioned in Section 2.1, this explanation method was used by most previous studies in automated medical coding Mullenbach et al. (2018); Kim et al. (2022); Dong et al. (2021); Teng et al. (2020); Cheng et al. (2023).
@@
The attention matrix in the cross-attention layer extracts information from the contextualized token representations encoded by RoBERTa (see Figure 5). The token representations are not guaranteed to be aligned with the input tokens. A token representation at position could represent any and multiple tokens in the document. Attention rollout considers all the model’s attention layers to calculate the feature attributions Abnar and Zuidema (2020). First, the attention matrices in each layer are averages across the heads. Then, the identity matrix is added to each layer’s attention matrix to represent the skip connections. Finally, the attention rollout is calculated recursively using Equation 4.
@@
Integrated Gradients (IntGrad)
Integrated Gradients (IntGrad) assigns an attribution score to each input feature by computing the integral of the gradients along the straight line path from the baseline to the input Sundararajan et al. (2017). Similar to InputXGradient, we used the L2-norm of the output to get the final attribution scores.
@@
AttnGrad
AttnGrad multiplies the attention Attention with the gradient of the model’s output with respect to the attention weights Serrano and Smith (2019):
@@ E.2 Results from all feature attribution methods
In the main paper, we only presented the results of selected feature attribution methods because of space constraints. Here, we present the results for all the feature attribution methods: Attention, AttGrad, Attention Rollout (Rollout), InputXGrad, Integrated gradients (IntGrad), Deeplift, and AttInGrad. We compare these methods with a random baseline (Rand), which randomly generates attribution scores. We present the plausibility results in Table 7, and the faithfulness results in Table 8.
@@ E.5 Unsupervised comparison on bigger test set
We present the plausibility results in Table 12. We observe that the results are similar to those of the main paper. However, the IGR produced substantially better attention-based explanations than in the main paper. In Figure 6, we inspect the inter-seed variance. We observe that IGR has no outliers. We, therefore, attribute the differences between this comparison and that in the main paper to none of the ten IGR runs happening to produce an outlier model. These results highlight the fragility of evaluating the attention-based feature attribution methods.
@@ Appendix F CO2 emissions
Appendix G Licenses
We used MIMIC-III version 1.4, which is distributed under a non-commercial license, as detailed here: https://physionet.org/content/mimiciii/view-license/1.4/. Consequently, all model weights released in this paper are also restricted to non-commercial use. However, the MDACE annotations and our code are available under the MIT License. We provide instructions to how to obtain the datasets in our GitHub repository.